# GPU-Accelerated Rule Evaluation and Evolution

### v1

@@ Abstract.
This paper introduces an innovative approach to boost the efficiency and scalability of Evolutionary Rule-based machine Learning (ERL), a key technique in explainable AI. While traditional ERL systems can distribute processes across multiple CPUs, fitness evaluation of candidate rules is a bottleneck, especially with large datasets. The method proposed in this paper, AERL (Accelerated ERL) solves this problem in two ways. First, by adopting GPU-optimized rule sets through a tensorized representation within the PyTorch framework, AERL mitigates the bottleneck and accelerates fitness evaluation significantly. Second, AERL takes further advantage of the GPUs by fine-tuning the rule coefficients via back-propagation, thereby improving search space exploration. Experimental evidence confirms that AERL search is faster and more effective, thus empowering explainable artificial intelligence.
@@ 1. Introduction
Evolutionary Rule-based Machine Learning (ERL) systems harness evolutionary algorithms to optimize rule sets in many machine learning domains, including supervised learning where fitness is evaluated based on accuracy on a pre-existing dataset (Miikkulainen, 2023; Shahrzad et al., 2022). Graphics Processing Units (GPUs) have long been used to accelerate such algorithms, including Mixed Integer Non-Linear Programming (MINLP), Central Force Optimization (CFO), Genetic Algorithms (GA), and Particle Swarm Optimization (PSO), with speedups ranging from 7x to 10,000x (Singh et al., 2014). Further enhancing such optimistic performance claims (Jähne, 2016), this paper introduces a novel tensor-based representation of logical rules utilizing PyTorch. This representation not only enhances evaluation speed but also fosters local optimization capability via back-propagation on GPUs, thereby yielding significant performance improvements. The approach is validated empirically through preliminary experiments in the UCI breast cancer dataset, leading to a number of future work directions.
@@ 2. Background
Evolving transparent explainable rule sets based on extended propositional logic is a viable approach to creating explainable AI systems. Contrary to traditional black-box models, stakeholders can easily understand and audit such systems (Shahrzad et al., 2022; Hemberg et al., 2014). However, there is untapped potential in enhancing these processes through GPU acceleration. Given that evolutionary algorithms are computationally intensive, requiring rapid, iterative rule evaluation, GPU acceleration should be effective. Yet it is rarely used, presumably because rule-sets do not naturally map to GPUs the same way s black-box models such as neural networks.
@@ 3. Method
The AERL approach transforms list-based propositional logic expressions (Shahrzad et al., 2022) into a computational model that can be efficiently processed using linear algebra and optimization techniques. By representing the rules as inequalities, AERL can process the logical conditions and their relationships in a structured numerical format. This conversion is pivotal because it allows the use of matrix multiplication to handle the evaluation of multiple rules simultaneously, thereby mapping the input features to the outcomes defined by the rules.

### v2

@@ Abstract.
This paper presents an extension to the EVOTER (Evolution of Transparent Explainable Rule Sets) framework, previously introduced to promote transparency and explainability in AI through the evolution of interpretable rule sets. While EVOTER demonstrated strong performance in generating compact and understandable rule-based models using a list-based grammar, the current work, Accelerated Evolutionary Rule-based Learning (AERL), builds upon this foundation by addressing scalability challenges. Specifically, AERL leverages GPU-accelerated computation and back-propagation fine-tuning within a PyTorch framework to significantly enhance the efficiency of rule evaluation and evolution. By introducing a tensorized numerical representation of rules and incorporating gradient-based optimization, AERL achieves superior computational speed and improved search-space exploration. Experimental results confirm the effectiveness of the approach in achieving faster convergence and better accuracy, offering a scalable path forward for transparent AI models.
@@ 1. Introduction
Evolutionary Rule-based Machine Learning (ERL) systems have been applied to a wide range of supervised learning tasks, where rule fitness is typically evaluated based on classification accuracy on a pre-existing dataset (Miikkulainen, 2023). A notable recent contribution in this space is EVOTER (Evolution of Transparent Explainable Rule Sets), which demonstrated how expressive rule sets can be evolved using a list-based grammatical framework (Shahrzad et al., 2024).
This paper presents an extension of EVOTER that addresses one of the primary bottlenecks in ERL systems: the computational cost of rule evaluation. Graphics Processing Units (GPUs) have long been used to accelerate optimization algorithms, including Mixed Integer Non-Linear Programming (MINLP), Central Force Optimization (CFO), Genetic Algorithms (GA), and Particle Swarm Optimization (PSO), with reported speedups ranging from 7x to 10,000x (Singh et al., 2014). Further enhancing such promising performance claims (Jähne, 2016), this work introduces a novel tensor-based representation of logical rules using the PyTorch framework. This representation enables massively parallel evaluation of rule sets on GPUs and supports local optimization through back-propagation, significantly accelerating both rule evaluation and search-space exploration.
By embedding rule sets into a numerical tensor format, the proposed method—Accelerated Evolutionary Rule-based Learning (AERL)—not only improves scalability but also leverages gradient-based fine-tuning to enhance evolutionary progress. Experimental validation on the UCI Breast Cancer dataset confirms the effectiveness of the approach, demonstrating substantial gains in both computational speed and model performance.
@@ 2. Background
Evolving transparent explainable rule sets based on extended propositional logic is a viable approach to creating explainable AI systems. Contrary to traditional black-box models, stakeholders can easily understand and audit such systems (Shahrzad et al., 2024; Hemberg et al., 2014). However, there is untapped potential in enhancing these processes through GPU acceleration. Given that evolutionary algorithms are computationally intensive, requiring rapid, iterative rule evaluation, GPU acceleration should be effective. Yet it is rarely used, presumably because rule-sets do not naturally map to GPUs the same way s black-box models such as neural networks.
@@ 3. Method
The AERL approach transforms list-based propositional logic expressions (Shahrzad et al., 2024) into a computational model that can be efficiently processed using linear algebra and optimization techniques. By representing the rules as inequalities, AERL can process the logical conditions and their relationships in a structured numerical format. This conversion is pivotal because it allows the use of matrix multiplication to handle the evaluation of multiple rules simultaneously, thereby mapping the input features to the outcomes defined by the rules.